OpenStudy (dancergirl45):

A greenhouse that specializes in growing geraniums is divided into sections. The number of geranium pots in each section depends on the number of sprinklers in that section. There are x(7x − 2) pots of red geraniums and (x + 3)2 pots of pink geraniums in a section with x sprinklers. Which simplified expression represents the total number of red and pink geranium pots in a section with x sprinklers?

OpenStudy (dancergirl45):

@phi

OpenStudy (dancergirl45):

I got 8x2 + 4x + 9 Is that wrong?

OpenStudy (dancergirl45):

@Ineedhelplz

OpenStudy (austinf):

What are your choices of answers? @Dancergirl45

OpenStudy (dancergirl45):

8x2 + 4x + 9 8x2 + 6x + 9 8x2 + 6x + 7 8x4 + 4x2 + 9

OpenStudy (dancergirl45):

I had A

OpenStudy (dancergirl45):

@phi am i correct?

OpenStudy (ineedhelplz):

I misread the answers, A is right...

OpenStudy (phi):

yes

OpenStudy (austinf):

x(7x-2)+(x+3)^2 The answer would be 8x^2+4x+9 A is correct.

OpenStudy (phi):

x(7x − 2) is 7x^2 - 2x (x + 3)^2 = x^2 + 6x + 9 ------------- 8x^2 +4x + 9

OpenStudy (dancergirl45):

Thank you and with this one (3x + 5)(5x − 1) Which of the following simplified expressions represents the area of the floor of the library building in square meters? 28x − 5 15x2 − 5 15x2 + 28x − 5 15x2 + 22x − 5

OpenStudy (dancergirl45):

I got D

OpenStudy (dancergirl45):

is that correct?

OpenStudy (austinf):

Yes, that is correct.

OpenStudy (cas_fangirl_14):

I found this http://openstudy.com/updates/4f3d8634e4b0cf389e29a16e

OpenStudy (dancergirl45):

@phi

OpenStudy (phi):

yes

OpenStudy (phi):

using FOIL: (3x + 5)(5x − 1) First: 3x*5x = 15*x*x = 15x^2 Outer: 3x*-1 = -3x Inner: 5*5x = 25x Last: 5*-1 = -5 15x^2 +22x - 5